# Condition Monitoring with Incomplete Data: An Integrated Variational Autoencoder and Distance Metric Framework

### v1

@@ I Introduction
Figure 1: The framework of the proposed method
@@ II-A Variational Autoencoder
A variational autoencoder is an autoencoder with regularized training to avoid overfitting and ensure that the latent space has good properties that enable a generative process. It maps the input to probabilistic distributions instead of a single point and can produce new data. At first, the input is decoded as a probabilistic distribution in latent space, and then a point is sampled from that distribution. After decoding the sample, reconstruction error is computed and propagated through the network. The latent space distribution is usually considered normal. The distributions generated by the decoder should closely resemble a standard normal distribution, ensuring that the latent space is properly regularized and facilitates the generation of new data. The variational autoencoder loss function is composed of a reconstruction term, which makes data reproduction possible with a minimum error like a conventional autoencoder, and a regularization term, which makes latent space distributions similar to standard normal. This loss function is expressed as the Kulback-Leibler divergence between the returned distribution and a standard Gaussian. Improving the regularization term and span of the whole latent space increases reconstruction error, so a trade-off should be considered in that case. The cost function of a variational autoencoder is described in (1), where $x$ is the input to the VAE, $f(z)$ is the output of the decoder,$q_{x}(z)$ is Gaussian distribution approximating $p(z|x)$ considering $z$ as the latent variable, and $p(z)$ is a standard Gaussian distribution.
@@
To improve the performance and robustness of the network, the coefficient $\beta$ is added to the base cost function of the autoencoder as shown in (2). A higher $\beta$ value increases the spread of data in the latent space, leading to better separation between data clusters. In a noisy environment, the noise will not have a major effect due to the distance between different groups of data, so the classification accuracy will be increased [15].
@@ II-B Distance Metrics
To establish a reliable health index for the system under different working conditions, we analyze the information from the latent space of the variational autoencoder. This approach leverages the VAE’s dimension reduction and feature extraction capabilities. The distance between each data point and the mean of the normal data in the latent space is calculated and used as the system’s health index. As the trained encoder of the VAE can map any data sample (even the data that has not been seen in the training) into a point in the latent space, the health index can be calculated for any data point. Different distance metrics are calculated and compared, including Euclidean distance, Manhattan distance, also known as city block, and Minkowski distance with the norm order of three, given by
$\begin{split}(p,q)=\sqrt{\sum_{i=1}^{n}(q_{i}-p_{i})^{2}}\end{split}$ (3)
$\begin{split}(p,q)={\sum_{i=1}^{n}|q_{i}-p_{i}|}\end{split}$ (4)
$\begin{split}(p,q)=({\sum_{i=1}^{n}|q_{i}-p_{i}|^{p}})^{\!1/p}\end{split}$ (5)
@@ III-B VAE Implementation Details
The designed VAE consists of a four-layer encoder and decoder with relu activation functions. The encoder in this architecture is designed as a fully connected Multilayer Perceptron (MLP) with layer sizes of 256, 128, 32, and 8 neurons in sequence. Correspondingly, the decoder reverses this configuration, following the neuron sequence of 8, 32, 128, and 256. The latent space is set to be five-dimensional, meaning that each input vector is encoded to a five-dimensional probabilistic distribution. The training is done over 500 epochs, with Adam optimizer with a learning rate of 5e-4. The $\beta$ in the VAE’s loss function is set to 20. Latent encodings of the testing phases are visualized in Fig.2, showing three of the five dimensions. The VAE successfully maps all the normal and degraded data to their respective distributions, and these classes are easily separable from each other within this latent space.
@@ III-C Health Index Assignment
The efficiency of an optimized VAE in mapping the data in a rich latent space is shown in the previous subsection. The normal, degraded and very faulty data are separable, and the level of the degradation of the system is related to the distance from the normal samples. The worse the condition of the system, the further from normal data. This assumption allows us to assign a health index to every new and unseen data sample. We define the health index as the distance between each representation of the data in the five-dimensional latent space from the mean of the normal training data.
@@
We evaluated the performance of three distance metrics (Euclidean, Manhattan, and Minkowski) for calculating the health index. The detailed comparison is presented in Table I. While all three metrics showed promising results, Euclidean distance demonstrated a slight advantage over the others.

### v2

@@ Abstract
©2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, including reprinting/republishing this material for advertising or promotional purposes, collecting new collected works for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
Figure 1: The framework of the proposed method1
@@ II-A Variational Autoencoder
A variational autoencoder is an autoencoder with regularized training to avoid overfitting and ensure that the latent space has good properties that enable a generative process. It maps the input to probabilistic distributions instead of a single point and can produce new data. At first, the input is encoded as a probabilistic distribution in latent space, and then a point is sampled from that distribution. After decoding the sample, reconstruction error is computed and propagated through the network. The distributions generated by the decoder should closely resemble a standard normal distribution, ensuring that the latent space is properly regularized and facilitates the generation of new data. The variational autoencoder loss function is composed of a reconstruction term, which makes data reproduction possible with a minimum error like a conventional autoencoder, and a regularization term, which makes latent space distributions similar to standard normal. This loss function is expressed as the Kulback-Leibler divergence between the returned distribution and a standard Gaussian. Improving the regularization term and span of the whole latent space increases reconstruction error, so a trade-off should be considered in that case. The loss function of a variational autoencoder is described in (1), where $x$ is the input to the VAE, $f(z)$ is the output of the decoder, $q_{x}(z)$ is Gaussian distribution approximating $p(z|x)$ considering $z$ as the latent variable, $p(z)$ is a standard Gaussian distribution, and $c$ is a constant balancing factor.
@@
To improve the performance and robustness of the network, the coefficient $\beta$ is added to the base loss function of the autoencoder as shown in (2). A higher $\beta$ value increases the spread of data in the latent space, leading to better separation between data clusters. In a noisy environment, the noise will not have a major effect due to the distance between different groups of data, so the classification accuracy will be increased [15].
@@ II-B Distance Metrics
To establish a reliable health index for the system under different working conditions, we analyze the information from the latent space of the variational autoencoder. This approach leverages the VAE’s dimension reduction and feature extraction capabilities. The distance between each data point and the mean of the normal data in the latent space is calculated and used as the system’s health index. As the trained encoder of the VAE can map any data sample (even the data that has not been seen in the training) into a point in the latent space, the health index can be calculated for any data point. Different distance metrics are calculated and compared, including Euclidean distance ($d_{E}$), Manhattan distance ($d_{T}$), also known as taxi-cab, and Minkowski distance ($d_{M}$) with the norm order of three, given by
$\begin{split}d_{E}(p,q)=\sqrt{\sum_{i=1}^{n}(q_{i}-p_{i})^{2}}\end{split}$ (3)
$\begin{split}d_{T}(p,q)={\sum_{i=1}^{n}|q_{i}-p_{i}|}\end{split}$ (4)
$\begin{split}d_{M}(p,q)=({\sum_{i=1}^{n}|q_{i}-p_{i}|^{p}})^{\!1/p}\end{split}$ (5)
@@
11footnotetext: This figure was prepared using images from www.flaticon.com.
@@ III-B VAE Implementation Details
The designed VAE consists of a four-layer encoder and decoder with relu activation functions. The encoder in this architecture is designed as a fully connected Multilayer Perceptron (MLP) with layer sizes of 256, 128, 32, and 8 neurons in sequence. Correspondingly, the decoder reverses this configuration, following the neuron sequence of 8, 32, 128, and 256. The latent space is set to be five-dimensional, meaning that each input vector is encoded to a five-dimensional probabilistic distribution. The training is done over 500 epochs, with Adam optimizer with a learning rate of 5e-4. The $\beta$ in the VAE’s loss function is set to 20, and $c$ is considered 1. Latent encodings of the testing phases are visualized in Fig.2, showing three of the five dimensions. The VAE successfully maps all the normal and degraded data to their respective distributions, and these classes are easily separable from each other within this latent space.
@@ III-C Health Index Assignment
The efficiency of an optimized VAE in mapping the data in a rich latent space is shown in the previous subsection. The normal, degraded and very faulty data are separable, and the level of the degradation of the system is related to the distance from the normal samples. The worse the condition of the system, the further from normal data. This assumption allows us to assign a health index to every new and unseen data sample. We define the health index as the distance between each encoded data point from the mean of the encoded normal condition data in the five-dimensional latent space.
@@
We evaluated the performance of three distance metrics (Euclidean, Manhattan, and Minkowski) for calculating the health index. The detailed comparison is presented in Table I. While all three metrics showed promising results, Euclidean distance demonstrated a slight advantage over the others. Therefore, it has been selected as the health index in this study.